x+3 over 6 = 2 over 3 + x-5 over 4
Answer by poliphob3.14(115)

Solution: (x+3)/6=2/3+(x-5)/4 Multiply both sides of equation with 12
2(x+3)=8+3(x-5); 2x+6=8+3x-15; x=13.
Proof:(13+3)/6=2/3+(13-5)/4; 16/6=2/3+8/4 8/3=2/3+2; 8/3=8/3.